Weekly LIPAc highlights (8/November-12/November)


Deuteron beam extraction for injector conditioning was carried out until 26 October and reached 30 mA @ 10% duty cycle. On 27 October proton beam extraction was started and reached quickly the current and duty level required to resume beam operation (total extracted current 20mA, duty 3%). 

RF system

One of the new Driver Anode Box (chain 1B) had an issue and was recovered by replacing it back to the old one (checks are on-going). Anode Boxes in the remaining chains will be replaced progressively in the next short term maintenances.

RF conditioning started on 28 October: in a half day, 1% duty (1ms, 10Hz) was reached. The system could very quickly ramp-up to 10% duty cycle, confirming the much improved stability of the overall RF system. Duty cycle reached 10% at 2ms, 50Hz. During last weekend at ~17% duty cycle there was an important RFQ vacuum interlock event and associated RFQ coupler temperature increase. The system recovered and was reconditioned at low duty cycle with no issue. A maximum of 10% duty cycle was kept with varying pulse length and repetition rate. Upon request by INFN, field flatness measurement were performed at several duty cycle (with fix pulse duration of 4ms) before going to the next step increasing the duty cycle up to 15%. Until winter break the plan is to have beam operation in day time and keep conditioning during the night time and weekend.

Beam Transport/BD

Before resuming beam operation, there was a very intense period of remote sessions with CIEMAT to solve some last minute issue and complete few remaining tasks. Power Supply (PS) OPI was modified to allow the use of the PS HPU04 which is perfectly functional but does not provide digital current readback (analog output are operative and processed through the OPI). 


In the remote sessions with CIEMAT the tasks for the cRIO SW update, the noise issues of BD accelerometer and the water temperature measurement at the D-slit were completed. BPM clock board was recovered by applying SW update.

Ancillaries/vacuum and cooling

Just before resuming the beam one vacuum gate valve (isolating a cryopump in the HEBT vacuum sector III) was blocked half open. Despite all the efforts it could not be recovered. A number of countermeasures have been taken to operate safely the vacuum system during the beam and the intervention on the valve for the next winter break is in preparation.

Improvements of the secondary cooling system (outdoor large water tank area) restarted to install protector against drop of ice and frozen snow.

The large Cryoplant 2nd storage buffer recently installed has been connected, conditioned and filled with some pure helium.

Control System

MPS test before beam operation was longer than planned and it was necessary to have CIEMAT to get all MPS signals for the DPlate diagnostic cleared. The MPS for the HEBT/BD was completed on 11 November in the morning.

Beam operation

A rehearsal for beam operation was done on 11 November and beam operation started on November 12.

Operations Supervisor and sub-systems operators observing traces of first beams in CCR
Remote participants observing traces of first beam on shared OPI

Disclaimer: the information reported is not meant to be technically complete and doesn’t cover all the activities currently carried out on LIPAc.